OpenStudy (anonymous):

Terry is going to purchase one item each from two quarter machines. The first quarter machine contains bracelets that are 4 different colors, including her favorite color, red. There are equal amounts of each color. The second quarter machine contains 1 sticky hand, 1 pencil topper, 2 erasers, 3 plastic rings, and 4 bouncing balls. What is the probability that Terry will get a red bracelet and a plastic ring by putting a quarter in each machine?

OpenStudy (the_fizicx99):

I skimmed the problem, >.< I'm getting off, but the probability formula I would use is \(\ \sf \dfrac{Favorable~Outcomes}{Total ~ number ~ of ~ outcomes} \)

OpenStudy (anonymous):

3/44?

OpenStudy (the_fizicx99):

\(\ \dfrac{1}{4} \) and \(\ \dfrac{3}{11} \) I may be wrong, feel free to correct me :)

OpenStudy (the_fizicx99):

Is 23/44 an option? Curious, tried a different approach

OpenStudy (anonymous):

no lol but what about 1/11?

OpenStudy (anonymous):

or 1/22?